# Nonsingular black holes and spherically symmetric objects in nonlinear electrodynamics with a scalar field

### v2

@@ Abstract
In general relativity with vector and scalar fields given by the Lagrangian ${\cal L}(F,\phi,X)$, where $F$ is a Maxwell term and $X$ is a kinetic term of the scalar field $\phi$, we study the linear stability of static and spherically symmetric objects without curvature singularities at their centers. We show that the background solutions are generally described by either purely electrically or magnetically charged objects with a nontrivial scalar-field profile. In theories with the Lagrangian ~$\tilde{{\cal L}}(F)+K(\phi,X)$, which correspond to nonlinear electrodynamics with a k-essence scalar field, angular Laplacian instabilities induced by vector-field perturbations exclude all the regular spherically symmetric solutions including nonsingular black holes. In theories described by the Lagrangian ${\cal L}=X+\mu(\phi)F^{n}$, where $\mu$ is a function of $\phi$ and $n$ is a constant, the absence of angular Laplacian instabilities of spherically symmetric objects requires that $n>1/2$, under which nonsingular black holes with apparent horizons are not present. However, for some particular ranges of $n$, there are horizonless compact objects with neither ghosts nor Laplacian instabilities in the small-scale limit. In theories given by ${\cal L}=X\kappa(F)$, where $\kappa$ is a function of $F$, regular spherically symmetric objects are prone to Laplacian instabilities either around the center or at spatial infinity. Thus, in our theoretical framework, we do not find any example of linearly stable nonsingular black holes.
@@ I Introduction
If we consider vector-tensor theories in nonlinear electrodynamics (NED) given by the Lagrangian ${\cal L}(F)$, where ${\cal L}$ is a nonlinear function of $F$, it is possible to realize nonsingular BHs without curvature singularities at $r=0$ [25, 26, 27, 28, 4, 7, 29, 30]. The NED Lagrangian accommodates Euler-Heisenberg theory [31] as well as Born-Infeld theory [32]. In such subclasses of NED theories, the resulting SSS BH solutions possess curvature singularities at $r=0$ [33, 34, 35, 36]. However, there are nonsingular electrically or magnetically charged BHs for some specific choices of the NED Lagrangian. Thus, the vector field with nonlinear Lagrangians of $F$ allows an interesting possibility for realizing regular BHs even at the classical level.
@@
After deriving conditions for the absence of ghosts and Laplacian instabilities of the five dynamical DOFs, we will apply them to three subclasses of ${\cal L}(F,\phi,X)$ theories: (i) ~$\tilde{\cal L}(F)+K(\phi,X)$, (ii) ${\cal L}=X+\mu(\phi)F^{n}$, and (iii) ${\cal L}=X\kappa(F)$. We will show that regular SSS objects realized by theories (i) and (iii), which include nonsingular BHs, are excluded by Laplacian instabilities of vector-field perturbations around the origin. In theories (ii), the absence of angular Laplacian instabilities requires the condition $n>1/2$, under which there are no regular BHs with apparent horizons. Thus, even by extending NED to more general theories with the Lagrangian ${\cal L}(F,\phi,X)$, we do not find even a single example of nonsingular BHs without instabilities.
@@ II Field equations on the SSS background
We consider theories in which the Lagrangian ${\cal L}$ in the matter sector depends on a scalar field $\phi$ and the two scalar products
@@
where $f$ and $h$ are functions of the radial distance $r$. For the positivity of $-g=(f/h)r^{4}\sin^{2}\theta$ in Eq. (2), we require that $f/h$ is positive. For the later convenience, we introduce the $r$-dependent function $N(r)$ satisfying
$f(r)=N(r)h(r)\,,$ (4)
@@
so that $A_{0}^{\prime}$ is known algebraically in terms of $h$, $N$, and its $r$ derivatives. Interestingly, the $\phi$-dependent terms completely vanish in Eq. (16).
@@ II.1 Electric case
which is finite at $r=0$. Substituting Eq. (21) into Eqs. (7) and (13), we know $F$ and ${\cal L}$ in terms of $h$, $N$, and their $r$ derivatives. Under the expansions (17) and (18) around $r=0$, it follows that
@@ III Perturbations on the SSS background
We choose the four gauge conditions $h_{t\theta}=0$, $h_{\theta\theta}=0$, $h_{\varphi\varphi}=0$, and $h_{\theta\varphi}=0$. In this case, the four components of $\xi^{\mu}$ under the infinitesimal coordinate transformation $x^{\mu}\to x^{\mu}+\xi^{\mu}$ are fixed. Then, the components of $h_{\mu\nu}$ are given by [50, 51, 52, 44]
@@
$h_{t\varphi}=-Q(t,r)(\sin\theta)Y_{l,\theta}(\theta),\;\;h_{rr}=f^{-1}(r)H_{2}(t,r)Y_{l}(\theta),$
@@ III.1 Second-order action
$\tilde{{\cal L}_{1}}$ $=$ ${\cal L}_{1}-s_{1}\biggl{[}\delta A_{0}^{\prime}-\dot{\delta A}_{1}$ (43)
@@
~$\tilde{{\cal L}}_{2}$ $=$ ${\cal L}_{2}-Lp_{1}\left[r\dot{W}-rQ^{\prime}+2Q-\frac{2{\cal L}_{,F}rA_{0}^{\prime}}{M_{\rm Pl}^{2}}\delta A-\chi\right]^{2}.$
@@ III.2 Stability of electric SSS objects
We first study the linear stability of electric SSS objects ($q_{M}=0$) by exploiting the action (42). This issue was addressed in Refs. [43, 44] as a special case of Maxwell-Horndeski theories, but we will revisit it to make a comparison with the stability of magnetic SSS objects. For electric objects, ~$\tilde{{\cal L}}_{1}$ is composed of even-parity perturbations alone, while ~$\tilde{{\cal L}}_{2}$ contains only odd-parity perturbations. In the following, we will study the case $h>0$, but we will also address the case $h<0$ at the end of this subsection.
@@ III.2.1 Odd-parity sector
We first consider perturbations in the odd-parity sector. Varying the Lagrangian ~$\tilde{{\cal L}}_{2}$ with respect to $Q$ and $W$, we obtain
@@
Substituting these expressions of $Q$ and $W$ and their $t$, $r$ derivatives into ~$\tilde{{\cal L}}_{2}$ and integrating it by parts, the second-order Lagrangian is expressed in the form
~$\tilde{{\cal L}}_{2}=\dot{\vec{\Psi}}_{\rm A}^{t}{\bm{K}}_{\rm A}\dot{\vec{\Psi}}_{\rm A}+\vec{\Psi}^{\prime t}{\bm{G}}_{\rm A}\vec{\Psi}_{\rm A}^{\prime}+\vec{\Psi}^{t}{\bm{M}}_{\rm A}\vec{\Psi}_{\rm A}\,,$ (47)
@@ III.2.2 Even-parity sector
In the even-parity sector, the product $H_{0}^{2}$ present in ${\cal L}_{1}$ disappears in ~$\tilde{\cal L}_{1}$ as a result of introducing the Lagrange multiplier $V$. Since ~$\tilde{\cal L}_{1}$ only contains terms linear in $H_{0}$, varying ~$\tilde{\cal L}_{1}$ with respect to $H_{0}$ puts constraints on other perturbed fields. We use this equation to express $h_{1}$ in terms of $\psi$, $V$, $\delta\phi$, and their derivatives. We also vary ~$\tilde{\cal L}_{1}$ with respect to $H_{1}$, $\delta A_{0}$, $\delta A_{1}$ and eliminate these fields from ~$\tilde{\cal L}_{1}$ by using their equations of motion. Up to boundary terms, we can express ~$\tilde{\cal L}_{1}$ in the following form
~$\tilde{{\cal L}}_{1}=\dot{\vec{\Psi}}_{\rm B}^{t}{\bm{K}}_{\rm B}\dot{\vec{\Psi}}_{\rm B}+\vec{\Psi}_{\rm B}^{\prime t}{\bm{G}}_{\rm B}\vec{\Psi}_{\rm B}^{\prime}+\vec{\Psi}_{\rm B}^{t}{\bm{M}}_{\rm B}\vec{\Psi}_{\rm B}+\vec{\Psi}_{\rm B}^{t}{\bm{Q}}_{\rm B}\vec{\Psi}_{\rm B}^{\prime}\,,$ (56)
@@
In theories where ${\cal L}$ is a function of $F$ alone, there are two dynamical perturbations $\psi$ and $V$ that propagate with the speed of light [42]. In current theories, we have an additional scalar perturbation $\delta\phi$ whose propagation speed is different from 1. To avoid the Laplacian instability of $\delta\phi$ along the radial direction, we require that $c_{r,\delta\phi}^{2}>0$.
@@
The above discussion is valid in the regime characterized by $h>0$, but we can perform a similar analysis for $h<0$. In the latter regime, the time-like and space-like properties of metric components $f~{}(=Nh)$ and $h$ are reversed compared to those in the former regime. The no-ghost conditions can be derived from the matrices ${\bm{G}}_{\rm A}$ and ${\bm{G}}_{\rm B}$ rather than ${\bm{K}}_{\rm A}$ and ${\bm{K}}_{\rm B}$. So long as ${\cal L}_{,F}>0$ and ${\cal L}_{,X}>0$, there are no ghosts in either odd-parity or even-parity sectors.
@@ III.3 Stability of magnetic SSS objects
where ~$\tilde{\cal L}_{\rm C}$ and ~$\tilde{\cal L}_{\rm D}$ are the Lagrangians containing perturbations in the sectors C and D, respectively. In the following, we will address the linear stability of magnetic SSS objects for the sectors C and D, in turn.
@@ III.3.1 Sector C
The Lagrangian ~$\tilde{{\cal L}}_{\rm C}$ is of the following form
~$\tilde{{\cal L}}_{\rm C}=\dot{\vec{\Psi}}_{\rm C}^{t}{\bm{K}}_{\rm C}\dot{\vec{\Psi}}_{\rm C}+\vec{\Psi}_{\rm C}^{\prime t}{\bm{G}}_{\rm C}\vec{\Psi}_{\rm C}^{\prime}+\vec{\Psi}_{\rm C}^{t}{\bm{M}}_{\rm C}\vec{\Psi}_{\rm C}+\vec{\Psi}_{\rm C}^{t}{\bm{Q}}_{\rm C}\vec{\Psi}_{\rm C}^{\prime}\,,$ (72)
@@
$(K_{\rm C})_{11}$ $=$ $\frac{M_{\rm Pl}^{2}L}{4hN^{3/2}(L-2)}>0\,,$ (73)
$(K_{\rm C})_{11}(K_{\rm C})_{22}-(K_{\rm C})_{12}^{2}$ $=$ $\frac{M_{\rm Pl}^{2}r^{4}{\cal L}_{,F}}{8h^{2}N^{3}(L-2)}>0\,.$ (74)
@@ III.3.2 Sector D
~$\tilde{{\cal L}}_{\rm D}=\dot{\vec{\Psi}}_{\rm D}^{t}{\bm{K}}_{\rm D}\dot{\vec{\Psi}}_{\rm D}+\vec{\Psi}_{\rm D}^{\prime t}{\bm{G}}_{\rm D}\vec{\Psi}_{\rm D}^{\prime}+\vec{\Psi}_{\rm D}^{t}{\bm{M}}_{\rm D}\vec{\Psi}_{\rm D}+\vec{\Psi}_{\rm D}^{t}{\bm{Q}}_{\rm D}\vec{\Psi}_{\rm D}^{\prime}\,,$ (78)
@@
If ${\cal L}$ contains nonlinear functions of $F$, the propagation speed of $\delta A$ is different from 1. The expression of $c_{\Omega,\delta A}^{2}$ coincides with the one derived in Ref. [42] for theories with ${\cal L}={\cal L}(F)$.
@@ IV.1 Electric case
$c_{\Omega,V}^{2}$ $=$ ~$\frac{\tilde{{\cal L}}_{,F}}{\tilde{{\cal L}}_{,F}+2F\tilde{{\cal L}}_{,FF}}\,,$ (91)
whereas all the other dynamical perturbations propagate with the speed of light. We can compute ~$\tilde{{\cal L}}_{,FF}$ by taking the $r$ derivatives of ~$\tilde{{\cal L}}_{,F}=q_{E}\sqrt{N}/(r^{2}A_{0}^{\prime})$ and $F=A_{0}^{\prime 2}/(2N)$, as ~$\tilde{{\cal L}}_{,FF}=\tilde{{\cal L}}_{,F}^{\prime}(r)/F^{\prime}(r)$. Then, Eq. (91) reduces to
@@
Since $A_{0}^{\prime}$ is given by Eq. (21), we can express $c_{\Omega,V}^{2}$ in terms of $h$, $N$, and their $r$ derivatives, as
@@ IV.2 Magnetic case
$c_{\Omega,\delta A}^{2}$ $=$ ~$\frac{\tilde{{\cal L}}_{,F}+2F\tilde{{\cal L}}_{,FF}}{\tilde{{\cal L}}_{,F}}\,,$ (97)
@@
~$\tilde{\cal L}(F)+K(\phi,X)=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,,$ (98)
with $F=-q_{M}^{2}/(2r^{4})$. Taking the $r$ derivative of Eq. (98) and exploiting Eqs. (9) and (11), we find that ~$\tilde{\cal L}_{,F}$ is written as
~$\tilde{\cal L}_{,F}$ $=$ $M_{\rm Pl}^{2}r^{2}[2N^{2}(r^{2}h^{\prime\prime}-2h+2)+2rhN(rN^{\prime\prime}+N^{\prime})$ (99)
@@
Differentiating this equation with respect to $r$, we can express ~$\tilde{\cal L}_{,FF}$ in terms of $N$, $h$, and their $r$ derivatives. Then, it follows that $c_{\Omega,\delta A}^{2}$ is completely identical to $c_{\Omega,V}^{2}$ for the electric configuration given by Eq. (93).
Using the expansions of $h$ and $N$ around $r=0$, the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is $-1$ and hence the odd-parity vector-field perturbation $\delta A$ is subject to angular Laplacian instability. This leads to the enhancement of the gravitational perturbation $\psi$. Then, the nonsingular magnetic SSS object cannot be present as a stable configuration.

### v3

@@ Abstract
In general relativity with vector and scalar fields given by the Lagrangian ${\cal L}(F,\phi,X)$, where $F$ is a Maxwell term and $X$ is a kinetic term of the scalar field $\phi$, we study the linear stability of static and spherically symmetric objects without curvature singularities at their centers. We show that the background solutions are generally described by either purely electrically or magnetically charged objects with a nontrivial scalar-field profile. In theories with the Lagrangian $\tilde{{\cal L}}(F)+K(\phi,X)$, which correspond to nonlinear electrodynamics with a k-essence scalar field, angular Laplacian instabilities induced by vector-field perturbations exclude all the regular spherically symmetric solutions including nonsingular black holes. In theories described by the Lagrangian ${\cal L}=X+\mu(\phi)F^{n}$, where $\mu$ is a function of $\phi$ and $n$ is a constant, the absence of angular Laplacian instabilities of spherically symmetric objects requires that $n>1/2$, under which nonsingular black holes with apparent horizons are not present. However, for some particular ranges of $n$, there are horizonless compact objects with neither ghosts nor Laplacian instabilities in the small-scale limit. In theories given by ${\cal L}=X\kappa(F)$, where $\kappa$ is a function of $F$, regular spherically symmetric objects are prone to Laplacian instabilities either around the center or at spatial infinity. Thus, in our theoretical framework, we do not find any example of linearly stable nonsingular black holes.
@@ I Introduction
[…]
@@
After deriving conditions for the absence of ghosts and Laplacian instabilities of the five dynamical DOFs, we will apply them to three subclasses of ${\cal L}(F,\phi,X)$ theories: (i) $\tilde{\cal L}(F)+K(\phi,X)$, (ii) ${\cal L}=X+\mu(\phi)F^{n}$, and (iii) ${\cal L}=X\kappa(F)$. We will show that regular SSS objects realized by theories (i) and (iii), which include nonsingular BHs, are excluded by Laplacian instabilities of vector-field perturbations around the origin. In theories (ii), the absence of angular Laplacian instabilities requires the condition $n>1/2$, under which there are no regular BHs with apparent horizons. Thus, even by extending NED to more general theories with the Lagrangian ${\cal L}(F,\phi,X)$, we do not find even a single example of nonsingular BHs without instabilities.
@@ II Field equations on the SSS background
[…]
@@
where $f$ and $h$ are functions of the radial distance $r$. For the positivity of $g=(f/h)r^{4}\sin^{2}\theta$ in Eq. (2), we require that $f/h$ is positive. For the later convenience, we introduce the $r$-dependent function $N(r)$ satisfying
f(r)=N(r)h(r),f(r)=N(r)h(r)\,, (4)
@@
so that $A_{0}^{\prime}$ is known algebraically in terms of $h$, NN, and its $r$ derivatives. Interestingly, the $\phi$-dependent terms completely vanish in Eq. (16).
@@ II.1 Electric case
[…]
@@ III Perturbations on the SSS background
We choose the four gauge conditions $h_{t\theta}=0$, $h_{\theta\theta}=0$, $h_{\varphi\varphi}=0$, and $h_{\theta\varphi}=0$. In this case, the four components of $\xi^{\mu}$ under the infinitesimal coordinate transformation $x^{\mu}\to x^{\mu}+\xi^{\mu}$ are fixed. Then, the components of $h_{\mu\nu}$ are given by [50, 51, 52, 44] [^2] In the version published in Physical Review D 111, 064051 (2025), there is a typo in the $h_{rr}$ component of Eq. (28). Here, we correct $f^{-1}(r)$ to $h^{-1}(r)$.
@@
$h_{t\varphi}=-Q(t,r)(\sin\theta)Y_{l,\theta}(\theta),\;\;h_{rr}=h^{-1}(r)H_{2}(t,r)Y_{l}(\theta),$
@@ III.1 Second-order action
$\hskip-19.91684pt\tilde{{\cal L}_{1}}$ $=$ ${\cal L}_{1}-s_{1}\biggl[\delta A_{0}^{\prime}-\dot{\delta A}_{1}$ (43)
@@
$\hskip-19.91684pt\tilde{{\cal L}}_{2}$ $=$ ${\cal L}_{2}-Lp_{1}\left[r\dot{W}-rQ^{\prime}+2Q-\frac{2{\cal L}_{,F}rA_{0}^{\prime}}{M_{\rm Pl}^{2}}\delta A-\chi\right]^{2}.$
@@ III.2 Stability of electric SSS objects
We first study the linear stability of electric SSS objects ($q_{M}=0$) by exploiting the action (42). This issue was addressed in Refs. [43, 44] as a special case of Maxwell-Horndeski theories, but we will revisit it to make a comparison with the stability of magnetic SSS objects. For electric objects, $\tilde{{\cal L}}_{1}$ is composed of even-parity perturbations alone, while $\tilde{{\cal L}}_{2}$ contains only odd-parity perturbations. In the following, we will study the case $h>0$, but we will also address the case $h<0$ at the end of this subsection.
@@ III.2.1 Odd-parity sector
We first consider perturbations in the odd-parity sector. Varying the Lagrangian $\tilde{{\cal L}}_{2}$ with respect to $Q$ and $W$, we obtain
@@
Substituting these expressions of $Q$ and $W$ and their $t$, $r$ derivatives into $\tilde{{\cal L}}_{2}$ and integrating it by parts, the second-order Lagrangian is expressed in the form
$\tilde{{\cal L}}_{2}=\dot{\vec{\Psi}}_{\rm A}^{t}{\bm{K}}_{\rm A}\dot{\vec{\Psi}}_{\rm A}+\vec{\Psi}^{\prime t}{\bm{G}}_{\rm A}\vec{\Psi}_{\rm A}^{\prime}+\vec{\Psi}^{t}{\bm{M}}_{\rm A}\vec{\Psi}_{\rm A}\,$, (47)
@@ III.2.2 Even-parity sector
In the even-parity sector, the product $H_{0}^{2}$ present in ${\cal L}_{1}$ disappears in $\tilde{\cal L}_{1}$ as a result of introducing the Lagrange multiplier $V$. Since $\tilde{\cal L}_{1}$ only contains terms linear in $H_{0}$, varying $\tilde{\cal L}_{1}$ with respect to $H_{0}$ puts constraints on other perturbed fields. We use this equation to express $h_{1}$ in terms of $\psi$, $V$, $\delta\phi$, and their derivatives. We also vary $\tilde{\cal L}_{1}$ with respect to $H_{1}$, $\delta A_{0}$, $\delta A_{1}$ and eliminate these fields from $\tilde{\cal L}_{1}$ by using their equations of motion. Up to boundary terms, we can express $\tilde{\cal L}_{1}$ in the following form
$\tilde{{\cal L}}_{1}=\dot{\vec{\Psi}}_{\rm B}^{t}{\bm{K}}_{\rm B}\dot{\vec{\Psi}}_{\rm B}+\vec{\Psi}_{\rm B}^{\prime t}{\bm{G}}_{\rm B}\vec{\Psi}_{\rm B}^{\prime}+\vec{\Psi}_{\rm B}^{t}{\bm{M}}_{\rm B}\vec{\Psi}_{\rm B}+\vec{\Psi}_{\rm B}^{t}{\bm{Q}}_{\rm B}\vec{\Psi}_{\rm B}^{\prime}\,$, (56)
@@
[…]
@@
The above discussion is valid in the regime characterized by $h>0$, but we can perform a similar analysis for $h<0$. In the latter regime, the time-like and space-like properties of metric components f(=Nh)f~(=Nh) and $h$ are reversed compared to those in the former regime. The no-ghost conditions can be derived from the matrices ${\bm{G}}_{\rm A}$ and ${\bm{G}}_{\rm B}$ rather than ${\bm{K}}_{\rm A}$ and ${\bm{K}}_{\rm B}$. So long as ${\cal L}_{,F}>0$ and ${\cal L}_{,X}>0$, there are no ghosts in either odd-parity or even-parity sectors.
@@ III.3 Stability of magnetic SSS objects
where $\tilde{\cal L}_{\rm C}$ and $\tilde{\cal L}_{\rm D}$ are the Lagrangians containing perturbations in the sectors C and D, respectively. In the following, we will address the linear stability of magnetic SSS objects for the sectors C and D, in turn.
@@ III.3.1 Sector C
The Lagrangian $\tilde{{\cal L}}_{\rm C}$ is of the following form
$\tilde{{\cal L}}_{\rm C}=\dot{\vec{\Psi}}_{\rm C}^{t}{\bm{K}}_{\rm C}\dot{\vec{\Psi}}_{\rm C}+\vec{\Psi}_{\rm C}^{\prime t}{\bm{G}}_{\rm C}\vec{\Psi}_{\rm C}^{\prime}+\vec{\Psi}_{\rm C}^{t}{\bm{M}}_{\rm C}\vec{\Psi}_{\rm C}+\vec{\Psi}_{\rm C}^{t}{\bm{Q}}_{\rm C}\vec{\Psi}_{\rm C}^{\prime}\,$, (72)
@@
$\hskip-14.22636pt(K_{\rm C})_{11}$ $=$ $\frac{M_{\rm Pl}^{2}L}{4hN^{3/2}(L-2)}>0\,$, (73)
$\hskip-14.22636pt(K_{\rm C})_{11}(K_{\rm C})_{22}-(K_{\rm C})_{12}^{2}$ $=$ $\frac{M_{\rm Pl}^{2}r^{4}{\cal L}_{,F}}{8h^{2}N^{3}(L-2)}>0\,.$ (74)
@@ III.3.2 Sector D
$\tilde{{\cal L}}_{\rm D}=\dot{\vec{\Psi}}_{\rm D}^{t}{\bm{K}}_{\rm D}\dot{\vec{\Psi}}_{\rm D}+\vec{\Psi}_{\rm D}^{\prime t}{\bm{G}}_{\rm D}\vec{\Psi}_{\rm D}^{\prime}+\vec{\Psi}_{\rm D}^{t}{\bm{M}}_{\rm D}\vec{\Psi}_{\rm D}+\vec{\Psi}_{\rm D}^{t}{\bm{Q}}_{\rm D}\vec{\Psi}_{\rm D}^{\prime}\,$, (78)
@@
[…]
@@ IV.1 Electric case
$c_{\Omega,V}^{2}$ $=$ $\frac{\tilde{{\cal L}}_{,F}}{\tilde{{\cal L}}_{,F}+2F\tilde{{\cal L}}_{,FF}}\,$, (91)
whereas all the other dynamical perturbations propagate with the speed of light. We can compute $\tilde{{\cal L}}_{,FF}$ by taking the $r$ derivatives of $\tilde{{\cal L}}_{,F}=q_{E}\sqrt{N}/(r^{2}A_{0}^{\prime})$ and $F=A_{0}^{\prime 2}/(2N)$, as $\tilde{{\cal L}}_{,FF}=\tilde{{\cal L}}_{,F}^{\prime}(r)/F^{\prime}(r)$. Then, Eq. (91) reduces to
@@
Since $A_{0}^{\prime}$ is given by Eq. (21), we can express $c_{\Omega,V}^{2}$ in terms of $h$, NN, and their $r$ derivatives, as
@@ IV.2 Magnetic case
$c_{\Omega,\delta A}^{2}$ $=$ $\frac{\tilde{{\cal L}}_{,F}+2F\tilde{{\cal L}}_{,FF}}{\tilde{{\cal L}}_{,F}}\,$, (97)
@@
$\tilde{\cal L}(F)+K(\phi,X)=\frac{M_{\rm Pl}^{2}}{r^{2}}\left(rh^{\prime}+h-1\right)\,$, (98)
with $F=-q_{M}^{2}/(2r^{4})$. Taking the $r$ derivative of Eq. (98) and exploiting Eqs. (9) and (11), we find that $\tilde{\cal L}_{,F}$ is written as
$\tilde{\cal L}_{,F}$ $=$ $M_{\rm Pl}^{2}r^{2}[2N^{2}(r^{2}h^{\prime\prime}-2h+2)+2rhN(rN^{\prime\prime}+N^{\prime})$ (99)
@@
Differentiating this equation with respect to $r$, we can express $\tilde{\cal L}_{,FF}$ in terms of $N$, $h$, and their $r$ derivatives. Then, it follows that $c_{\Omega,\delta A}^{2}$ is completely identical to $c_{\Omega,V}^{2}$ for the electric configuration given by Eq. (93).
Using the expansions of $h$ and $N$ around $r=0$, the leading-order contribution to $c_{\Omega,\delta A}^{2}$ is 1-1 and hence the odd-parity vector-field perturbation $\delta A$ is subject to angular Laplacian instability. This leads to the enhancement of the gravitational perturbation $\psi$. Then, the nonsingular magnetic SSS object cannot be present as a stable configuration.